Problem 10.[10 pts]. Suppose Bank of America stock has a beta of 1.65, whereas Boeing stock has a beta a- b- c- d- of 1.40. If the risk-free interest rate is 4% and the expected return of the market portfolio is 12%, according to the CAPM, (Show your calculations) What is the expected return of Bank of America stock? What is the expected return of Boeing stock? What is the beta of a portfolio that consists of 60% Bank ofAmerica stock and 40% Boeing stock? What is the expected return of a portfolio that consists of 60% of Bank of America and 40% Boeing stock? (Make sure to solve it in two ways).
Answer #1

Problem 10 :

a. Bank of America has a beta of 1.65

So, the return according to the CAPM model is :

Re = Rf + beta( Rm - Rf)

where Rf is the risk free rate

Rm is the return on the market

Re = 4 + 1.65 ( 12 - 4)

= 17.2%

b. The expected return on Boeing stock is :

using the same formula for CAPM , the expected return is :

Re = 4 + 1.4 ( 12 - 4)

= 4 + 11.2

=15.2%

c. Beta of a portfolio is :

weight of Bank of America stock *Beta of the stock + weight of Boeing stock *beta of Boeing stock

= (0.6) *1.65 + 0.4*1.4

=0.99 + 0.56

=1.55

d. the expected return of the portfolio:

weight of Bank of America* expected return + weight of Boeing stock* expected return on the stock

= 0.6*17.2 + 0.4*15.2

=10.32 + 6.08

= 16.4%